OpenStudy (anonymous):

Which of the following functions is continuous at x = 3?

OpenStudy (anonymous):

A. B. C. D

OpenStudy (anonymous):

D.all are continuos at x = 3

OpenStudy (michele_laino):

hint: the subsequent function: \[\Large f\left( x \right) = \frac{{{x^2} - 9}}{{x + 3}}\] is continuous at the subsequent set: \[\Large \left\{ {x \in {\mathbf{R}},\quad x \ne - 3} \right\}\]

OpenStudy (haseeb96):

A is correct answer

OpenStudy (anonymous):

isint c right too though?

OpenStudy (haseeb96):

no

OpenStudy (usukidoll):

Ah. I thought I misunderstood this. nevermind XD But yeah don't pick functions that are fractions because it causes restrictions and discontinuity.

OpenStudy (usukidoll):

I'll demonstrate why A is correct...  use difference of square formula for the numerator